### v1

# mini-vec2vec: Scaling Universal Geometry Alignment with Linear Transformations

Abstract
We build upon vec2vec, a procedure designed to align text embedding spaces without parallel data. vec2vec finds a near-perfect alignment, but it is expensive and unstable. We present mini-vec2vec, a simple and efficient alternative that requires substantially lower computational cost and is highly robust. Moreover, the learned mapping is a linear transformation. Our method consists of three main stages: a tentative matching of pseudo-parallel embedding vectors, transformation fitting, and iterative refinement. Our linear alternative exceeds the original instantiation of vec2vec by orders of magnitude in efficiency, while matching or exceeding their results. The method’s stability and interpretable algorithmic steps facilitate scaling and unlock new opportunities for adoption in new domains and fields.
Algorithm 1 Unsupervised Embedding Alignment
1:Embedding matrices $\mathbf{X}_{A}\in\mathbb{R}^{n_{A}\times d}$, $\mathbf{X}_{B}\in\mathbb{R}^{n_{B}\times d}$
2:Number of clusters $c$, neighbors $k$, iterations $T$, smoothing parameter $\alpha$, runs $s$, clusters $c^{\prime}$, neighbors $k^{\prime}$, sample size $n_{s}$
3:Linear transformation matrix $\mathbf{W}$
4:Normalize embeddings: $\hat{\mathbf{X}}_{A},\hat{\mathbf{X}}_{B}\leftarrow\text{center and normalize to unit sphere}$
5:Create pseudo-pairs: $\mathcal{C}\leftarrow\operatorname{AnchorAlignment}(\hat{\mathbf{X}}_{A},\hat{\mathbf{X}}_{B},c,k,s)$ $\triangleright$ See Alg. 2
6:Estimate initial mapping: $\mathbf{W}\leftarrow\text{Procrustes analysis on }\mathcal{C}$
7:Refine-1: Apply matching-based refinement on $\mathbf{W}$ $\triangleright$ See Alg. 3
8:Refine-2: Apply clustering-based refinement on $\mathbf{W}$ $\triangleright$ See Alg. 4
9:return $\mathbf{W}$
1 Introduction
Representation learning has revolutionized natural language processing by producing embedding models that capture rich semantic structure in high-dimensional vector spaces. These learned representations encode complex linguistic relationships, enabling downstream tasks such as semantic similarity and information retrieval. However, different models, training procedures, or even different random initializations of the same architecture can produce embedding spaces that, while semantically meaningful, are not directly comparable due to arbitrary rotations, reflections, and translations in their coordinate systems.
The problem of aligning embedding spaces without parallel supervision – that is, without access to pairs of embeddings that represent identical semantic content – has emerged as a fundamental challenge with far-reaching implications. Successful unsupervised alignment has multiple implications across many domains: multilingual natural language processing (Conneau et al., 2018; Zhang et al., 2017), cross-modal alignment between text and image representations (Maniparambil et al., 2024), and privacy concerns (Song and Raghunathan, 2020; Morris et al., 2023).
The recent vec2vec method (Jha et al., 2025) demonstrated that unsupervised alignment is indeed possible using CycleGAN (Zhu et al., 2020), an adversarial training framework. Their approach is theoretically grounded in the Platonic Representation Hypothesis (Huh et al., 2024), which posits that well-trained representations, regardless of their specific architectures or training procedures, tend to converge toward geometrically similar spaces. Under this hypothesis, semantic relationships are preserved across models in terms of relative distances and angles, suggesting that alignment should be achievable through geometric transformations.
While the vec2vec approach has proven effective, adversarial training frameworks bring inherent challenges that limit their practical applicability. GANs are notoriously computationally intensive, requiring careful hyperparameter tuning and often exhibiting training instability characterized by mode collapse, oscillatory behavior, or failure to converge (Saxena and Cao, 2023; Goodfellow, 2017). These methods typically demand substantial computational resources, including GPU acceleration and large amounts of training data – factors that may prohibit their utilization by most researchers and users.
Our point of departure is still universal geometry, but we pursue a different approach. We draw inspiration from the remarkable success of supervised alignment approaches, particularly neural network stitching methods (Lenc and Vedaldi, 2015; Bansal et al., 2021), which have demonstrated that simple affine transformations often suffice to align representations. Unfortunately, stitching requires access to pairs of matching representations. Our key methodological insight is that with universal geometry, structural similarities should be detectable and exploitable, even without adversarial training. Specifically, the relative arrangement of semantic clusters, the preservation of neighborhood structures, and the overall geometric organization of concepts should provide sufficient signal for alignment, even in the complete absence of parallel supervision. To this end, we utilize the Relative Representation framework of Moschella et al. (2023), which allows us to operate in a universal space agnostic of arbitrary coordinate systems.
Our contribution is mini-vec2vec, a simple and robust pipeline with competitive alignment performance offering several key advantages over adversarial approaches: computational efficiency (running on commodity CPU hardware), training stability (deterministic components with controllable stochastic elements), interpretability (each step has a clear geometric interpretation), and sample efficiency (effective with substantially fewer training examples). Through comprehensive experiments, we demonstrate that this simple approach not only matches but often exceeds the performance of the more complex adversarial method while requiring orders of magnitude less computational resources.
2 Preliminaries
The hypothesis that neural networks converge to geometrically similar representations has been explored across multiple domains and scales. Early work by Olah (2015) revealed striking structural regularities in representations of different convolutional networks, showing distances between object representations are roughly the same regardless of architecture. These findings suggested that the geometric organization of learned features might be more universal than previously assumed. The Platonic Representation Hypothesis, formally articulated by Huh et al. (2024), represents the most comprehensive theoretical framework for understanding universal geometry. The hypothesis posits that well-trained representations converge toward a shared geometric structure that reflects the underlying structure of the data domain.
Moschella et al. (2023), inspired by Olah (2015), utilized these observations by proposing relative representations as a mathematically principled framework for comparing embedding spaces. Their key insight was that while absolute coordinates may vary arbitrarily across models, relative similarity structures – captured through pairwise distances to a shared set of anchor data points – remain remarkably stable. This work provided both theoretical foundation and empirical evidence for the existence of model-agnostic geometric properties.
Given a set of anchor points $\{\mathbf{a}_{i}\}_{i=1}^{k}$ in an embedding space, the relative representation of any point $\mathbf{x}$ is defined as:
$\mathbf{r}(\mathbf{x})=[f(\mathbf{x},\mathbf{a}_{1}),f(\mathbf{x},\mathbf{a}_{2}),\ldots,f(\mathbf{x},\mathbf{a}_{k})]^{T},$
where $f(\cdot,\cdot)$ is a similarity function (typically cosine similarity). Two representations from different models will have close relative representations if they represent similar objects.
Algorithm 2 Algorithm for creating pseudo-pairs using noisy anchors
1:Normalized embeddings $\hat{\mathbf{X}}_{A},\hat{\mathbf{X}}_{B}$, clusters $c$, neighbors $k$, runs $s$
2:Set of pseudo-parallel pairs $\mathcal{C}$
3:Initialize collections: $\mathcal{R}_{A}\leftarrow[],\mathcal{R}_{B}\leftarrow[]$
4:for $i=1$ to $s$ do $\triangleright$ Multiple runs for robustness
5:  Cluster both spaces: $\{\mathbf{c}^{A}_{i,j}\}\leftarrow\text{K-means}(\hat{\mathbf{X}}_{A})$, $\{\mathbf{c}^{B}_{i,j}\}\leftarrow\text{K-means}(\hat{\mathbf{X}}_{B})$
6:  Compute similarities: $\mathbf{S}^{A},\mathbf{S}^{B}\leftarrow\text{cosine similarities between centroids}$
7:  Find correspondence: $\mathbf{P}^{*}_{i}\leftarrow\text{solve QAP}(\mathbf{S}^{A},\mathbf{S}^{B})$ using multiple 2-OPT runs
8:  Build relative representations: $\mathbf{r}^{A}_{i},\mathbf{r}^{B}_{i}\leftarrow\text{cosine similarities to anchors}$
9:  Store: append $\mathbf{r}^{A}_{i},\mathbf{r}^{B}_{i}$ to $\mathcal{R}_{A},\mathcal{R}_{B}$
10:end for
11:Concatenate: $\mathbf{r}^{A}\leftarrow\text{concatenate all runs in }\mathcal{R}_{A}$
12:Concatenate: $\mathbf{r}^{B}\leftarrow\text{concatenate all runs in }\mathcal{R}_{B}$
[…]
14:Create pseudo-pairs: $\mathcal{C}\leftarrow\left\{(\mathbf{x}^{A}_{i},\frac{1}{k}\sum_{j}\mathbf{x}^{B}_{j}):\mathbf{r}^{B}_{j}\in\mathcal{N}_{k}(\mathbf{r}^{A}_{i})\right\}$
15:return $\mathcal{C}$
3 Problem Formulation
We follow the setup laid out in vec2vec. We assume access to two pools of text embeddings, each from a different encoder, and our task is to align their embedding spaces without paired examples. Similar to vec2vec, we learn a parameterized mapping (neural network or linear transformation) rather than a one-to-one matching between data points. This approach offers key advantages:
  • Generalization: New embeddings can be translated without re-applying the algorithm
  • Robustness: Works when one-to-one point matching doesn’t exist, requiring only similar data distributions
  • Out-of-distribution capability: The parameterized nature enables generalization beyond the training data
The universal geometry hypothesis indicates that such a task can be possible: “good enough” models learn equivalent geometries where cosine similarities remain approximately invariant across embedding spaces. These geometric similarities enable the identification of this alignment.
Let $\mathcal{X}_{A}=\{\mathbf{x}^{A}_{i}\}_{i=1}^{n_{A}}$ and $\mathcal{X}_{B}=\{\mathbf{x}^{B}_{j}\}_{j=1}^{n_{B}}$ denote two sets of $d$-dimensional embeddings, where $\mathbf{x}^{A}_{i},\mathbf{x}^{B}_{j}\in\mathbb{R}^{d}$. These embeddings represent semantic content from potentially different models, architectures, or training procedures, but we assume they encode similar semantic information from the same underlying distribution $\mathcal{D}$. Unlike more traditional setups, there does not exist a pairing between data points, i.e., there is no overlap between the sentences themselves.
We organize these embeddings into matrices $\mathbf{X}_{A}\in\mathbb{R}^{n_{A}\times d}$ and $\mathbf{X}_{B}\in\mathbb{R}^{n_{B}\times d}$, where each row corresponds to a single embedding vector. The fundamental challenge is that while both spaces may contain similar semantic structures, they are related by an unknown transformation. Our goal is to learn a transformation function $f:\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that $f(\mathbf{x}^{A}_{i})\approx\mathbf{x}^{B}_{j}$ when $\mathbf{x}^{A}_{i}$ and $\mathbf{x}^{B}_{j}$ represent similar semantic content. [^1] More generally, the setup allows for learning a pair of functions $f_{1},f_{2}$ sending embeddings to a third space, such that $f_{1}(\mathbf{x}^{A}_{i})\approx f_{2}(\mathbf{x}^{B}_{j})$. We will not take advantage of this flexibility in the paper.
Algorithm 3 Refine-1: Matching-Based Refinement
1:Embedding matrices $\hat{\mathbf{X}}_{A},\hat{\mathbf{X}}_{B}$, initial transformation $\mathbf{W}$
2:Iterations $T$, smoothing parameter $\alpha$, neighbors $k^{\prime}$, sample size $n_{s}$
3:Refined transformation matrix $\mathbf{W}$ after matching-based refinement
4:for $t=1$ to $T$ do
5:  Sample embeddings: $\mathbf{X}_{\text{sample}}\leftarrow\text{random sample of }n_{s}\text{ rows from }\hat{\mathbf{X}}_{A}$
6:  Transform samples: $\mathbf{X}_{\text{transformed}}\leftarrow\mathbf{X}_{\text{sample}}\mathbf{W}$
[…]
8:  Average neighbors: $\mathbf{X}_{\text{matched}}\leftarrow\text{mean of neighbors for each sample}$
9:  Construct pseudo-pairs: $\mathcal{C}^{\prime}\leftarrow\{(\mathbf{X}_{\text{sample}},\mathbf{X}_{\text{matched}})\}$
10:  Estimate new mapping: $\mathbf{W}_{\text{new}}\leftarrow\text{Procrustes analysis on }\mathcal{C}^{\prime}$
11:  Update with smoothing: $\mathbf{W}\leftarrow(1-\alpha)\mathbf{W}+\alpha\mathbf{W}_{\text{new}}$
12:end for
Algorithm 4 Refine-2: Clustering-Based Refinement
1:Embedding matrices $\hat{\mathbf{X}}_{A},\hat{\mathbf{X}}_{B}$, initial transformation $\mathbf{W}$
2:Smoothing parameter $\alpha$, number of clusters $c^{\prime}$
3:Refined transformation matrix $\mathbf{W}$ after clustering-based refinement
4:Cluster space A: $\{\mathbf{c}^{A}_{j}\}\leftarrow\text{K-means}(\hat{\mathbf{X}}_{A},c^{\prime})$
5:Transform centroids: $\{\mathbf{c}^{A\to B}_{j}\}\leftarrow\{\mathbf{c}^{A}_{j}\mathbf{W}\}$
6:Cluster space B with seeds: $\{\mathbf{c}^{B}_{j}\}\leftarrow\text{K-means}(\hat{\mathbf{X}}_{B},c^{\prime},\text{init}=\{\mathbf{c}^{A\to B}_{j}\})$
7:Construct pairs $\mathcal{C}^{\prime}$ by matching cluster centroids
8:Estimate new mapping: $\mathbf{W}_{\text{new}}\leftarrow\text{Procrustes analysis on }\mathcal{C}^{\prime}$
9:Update with smoothing: $\mathbf{W}\leftarrow(1-\alpha)\mathbf{W}+\alpha\mathbf{W}_{\text{new}}$
10:return $\mathbf{W}$
vec2vec mini-vec2vec
$M_{1}$ $M_{2}$ Top-1 $\uparrow$ Rank $\downarrow$ Initial Refine-1 Refine-2 Top-1 $\uparrow$ Rank $\downarrow$
gtr 0.99 1.19 (0.1) 0.25 (0.02) 0.58 (0.00) 0.57 (0.00) 0.99 (0.00) 1.02 (0.00)
gran. stel. 0.98 1.05 (0.0) 0.38 (0.00) 0.59 (0.00) 0.60 (0.00) 0.99 (0.00) 1.03 (0.00)
e5 0.98 1.11 (0.0) 0.28 (0.00) 0.52 (0.00) 0.53 (0.00) 0.99 (0.00) 1.05 (0.00)
gran. 0.99 1.02 (0.0) 0.31 (0.02) 0.58 (0.00) 0.57 (0.00) 1.00 (0.00) 1.01 (0.00)
gtr stel. 0.99 1.03 (0.0) 0.29 (0.00) 0.53 (0.00) 0.54 (0.00) 0.98 (0.00) 1.03 (0.00)
e5 0.84 2.88 (0.2) 0.22 (0.04) 0.48 (0.00) 0.49 (0.00) 0.98 (0.00) 1.06 (0.00)
gran. 0.98 1.08 (0.0) 0.37 (0.02) 0.58 (0.00) 0.59 (0.00) 0.99 (0.00) 1.02 (0.00)
stel. gtr 1.00 1.10 (0.0) 0.28 (0.02) 0.54 (0.00) 0.54 (0.00) 0.96 (0.00) 1.10 (0.00)
e5 1.00 1.00 (0.0) 0.43 (0.00) 0.61 (0.00) 0.62 (0.00) 1.00 (0.00) 1.00 (0.00)
gran. 0.99 2.20 (0.2) 0.30 (0.00) 0.54 (0.00) 0.56 (0.00) 0.99 (0.00) 1.04 (0.00)
e5 gtr 0.82 2.56 (0.0) 0.22 (0.03) 0.49 (0.00) 0.51 (0.00) 0.96 (0.00) 1.10 (0.00)
stel. 1.00 1.00 (0.0) 0.44 (0.01) 0.64 (0.00) 0.65 (0.00) 1.00 (0.00) 1.00 (0.00)
Table 1: Results show mean $\pm$ standard deviation over 3 runs. Bold indicates best performance. Underline indicates nearly equally good performance ($\leq 0.01$ difference). Bold and underline combined indicate better performance, but only within a $0.01$ margin.
4 Methodology
4.1 Motivation
In this work, we restrict our attention to linear transformations, i.e., $f(\mathbf{x})=\mathbf{W}\mathbf{x}$ where $\mathbf{W}\in\mathbb{R}^{d\times d}$. Moreover, we will constrain our search space to orthogonal transformations, i.e., matrices $\mathbf{W}$ that satisfy $\mathbf{W}^{T}\mathbf{W}=\mathbf{I}$. This is motivated by the universal geometry hypothesis – we expect cosine similarities to be preserved across embedding spaces. Later on, we will relax this assumption. Ultimately, the matrix will be the exponentially-smoothed average of a sequence of orthogonal matrices, enforcing orthogonality softly.
Our linear approach addresses key limitations of existing methods: GANs are notoriously unstable and hard to train with many hyperparameters, require extensive computational resources and training time, and often need multiple seeds to achieve good solutions. In contrast, we know that in supervised cases (model stitching), there usually exists an affine mapping between latent spaces, suggesting the linear case is realizable – the question is how to find it efficiently. Our method requires significantly less data (60k vs. 1 million samples), runs on a CPU rather than requiring extensive GPU resources, and provides more stable training dynamics.
Our method resembles Hoshen and Wolf (2018)’s approach for word embeddings, proceeding in three main stages: approximate matching, transformation fitting, and iterative refinement. This scheme is very powerful, and we apply a similar method. We find that the approximate matching stage especially requires more delicate treatment in our case, so their approach is not applicable as-is.
The stages incrementally improve alignment quality – each stage builds upon structural correspondences to establish increasingly accurate alignments:
  • Approximate Matching: Find approximate pairs of embeddings from both spaces using landmark points as anchors for relative representations
  • Mapping Estimation: Learn a linear orthogonal mapping from approximate matchings, where linearity and orthogonality provide an inductive bias to smooth out incorrect pairings
  • Iterative Refinement: Use the initial solution as a stepping stone for better matching/mapping, iterating until convergence or budget exhaustion
The overall process is described in Algorithm 1.
4.2 Approximate Matching
Preprocessing.
We first address arbitrary translations and scalings across embedding spaces. Both spaces are centered around their respective means and normalized to the unit hypersphere:
$\boldsymbol{\mu}_{A}$ $=\frac{1}{n_{A}}\mathbf{X}_{A}^{T}\mathbf{1}_{n_{A}},\quad\tilde{\mathbf{X}}_{A}=\mathbf{X}_{A}-\mathbf{1}_{n_{A}}\boldsymbol{\mu}_{A}^{T}$
$\hat{\mathbf{X}}_{A}[i,:]$ $=\frac{\tilde{\mathbf{X}}_{A}[i,:]}{\|\tilde{\mathbf{X}}_{A}[i,:]\|_{2}},\quad\text{similarly for space }B$
where $\boldsymbol{\mu}_{A}$ is the mean of space $A$, $\mathbf{1}_{n_{A}}$ is the vector of ones of length $n_{A}$, and $\|\cdot\|_{2}$ denotes the L2 norm.
Centroid Matching.
To establish structural correspondences, we employ an anchor discovery procedure based on multiple runs of clustering and matching. Each run consists of K-means clustering followed by solving a quadratic assignment problem (QAP) for cluster correspondence.
In each run, we first perform K-means clustering on both embedding spaces independently to obtain cluster centroids $\{\mathbf{c}^{A}_{j}\}$ and $\{\mathbf{c}^{B}_{j}\}$ where $j=1,\ldots,k$. We then compute structural similarity matrices between cluster centroids:
$S^{A}_{ij}$ $=\cos(\mathbf{c}^{A}_{i},\mathbf{c}^{A}_{j}),\quad S^{B}_{ij}=\cos(\mathbf{c}^{B}_{i},\mathbf{c}^{B}_{j})$
where $\mathbf{c}^{A}_{i}$ and $\mathbf{c}^{B}_{i}$ are the $i$-th cluster centroids in spaces $A$ and $B$ respectively.
We find the optimal matching between cluster centroids by solving the quadratic assignment problem:
$\mathbf{P}^{*}=\operatorname{argmax}_{\mathbf{P}\in\Pi_{k}}\text{Tr}\left(\mathbf{S}_{A}^{T}\mathbf{P}\mathbf{S}_{B}\mathbf{P}^{T}\right)$
where $\mathbf{P}$ is a permutation matrix from the set $\Pi_{k}$ of all $k\times k$ permutation matrices. For solving QAP, we use $2$-OPT (running 30 times and choosing the one with the highest alignment score for further robustness).
Relative Representations.
Once we have identified an alignment between centroids, we use the aligned centroids as anchors. For each run within the anchor discovery phase, we represent each embedding through its relationships to anchor points. For embedding $\mathbf{x}^{A}_{i}$ in a given run, we construct:
$\mathbf{r}^{A}_{i}=[\cos(\mathbf{x}^{A}_{i},\mathbf{c}^{A}_{1}),\cos(\mathbf{x}^{A}_{i},\mathbf{c}^{A}_{2}),\ldots,\cos(\mathbf{x}^{A}_{i},\mathbf{c}^{A}_{k})]^{T}$
where $\mathbf{r}^{A}_{i}$ is the relative representation of the $i$-th embedding in space $A$ with respect to the $k$ cluster centroids from that run.
Robustification through Ensembling.
We improve robustness by concatenating relative representations from multiple runs:
$\mathbf{r}^{A}_{i}=[\mathbf{r}^{A}_{i,1};\mathbf{r}^{A}_{i,2};\ldots;\mathbf{r}^{A}_{i,s}]\in\mathbb{R}^{sk}$
This concatenation creates richer structural signatures that are more robust to clustering initialization randomness.
Anchor alignment algorithm is provided in Algorithm 2.
4.3 Mapping Estimation
We construct pseudo-parallel pairs by matching embeddings by similarity in relative space, averaging the $k$ nearest neighbors to reduce noise:
$\mathcal{C}=\left\{\left(\mathbf{x}^{A}_{i},\frac{1}{k}\sum_{\mathbf{r}^{B}_{j}\in\mathcal{N}_{k}(\mathbf{r}^{A}_{i})}\mathbf{x}^{B}_{j}\right):i=1,\ldots,n_{A}\right\}$
where $\mathcal{N}_{k}(\mathbf{r}^{A}_{i})$ denotes the $k$ nearest neighbors to $\mathbf{r}^{A}_{i}$ in the concatenated relative representation space. Note that the neighborhood information is obtained from the relative space (which is shared) and averaging takes place in the absolute space $B$. The orthogonal transformation is obtained via Procrustes analysis: $\mathbf{W}^{*}=\mathbf{V}\mathbf{U}^{T}$ where $\mathbf{U}\boldsymbol{\Sigma}\mathbf{V}^{T}=\text{SVD}(\mathbf{A}^{T}\mathbf{B})$, and $\mathbf{A}$, $\mathbf{B}$ are matrices formed from the pseudo-parallel pairs in $\mathcal{C}$.
4.4 Iterative Refinement
The initial transformation provides a coarse alignment that we refine with two complementary strategies. This refinement is crucial to improve alignment quality, with Refine-2 helping Refine-1 escape local minima. Pseudo-code is provided in Algorithm 3 & 4.
Refine-1: Matching-Based Refinement.
This refinement strategy proceeds by transforming embeddings from space $A$ to $B$ using $\mathbf{W}$, and averaging their nearest neighbors in space $B$ based on cosine similarity, and obtaining a new orthogonal transformation with Procrustes analysis – similar to the process described in the Mapping Estimation stage (Section 4.3), with the crucial difference that now the similarity is performed in the ambient space $B$ rather than relative space. The method is cheap and it runs for many iterations (50–100) using subsampling to boost computational efficiency.
We use exponential smoothing for transformation updates:
$\mathbf{W}^{(t+1)}=(1-\alpha)\mathbf{W}^{(t)}+\alpha\mathbf{W}_{\text{new}}^{(t)}$
with $\alpha=0.5$, though the results appear robust to other choices as well.
Refine-2: Clustering-Based Refinement.
This strategy aims to improve the large-scale matching of embeddings by matching cluster centroids. We begin by clustering the embeddings in space $A$. We then transform the $A$ cluster centroids, and cluster the $B$ embeddings with the transformed $A$ centroids as initialization. The use of cluster centroids creates more stable pseudo-parallel pairs based on cluster assignments rather than individual nearest neighbors. The use of the transformed centroids as initialization follows from the assumption that they cannot be far off from cluster centroids in space $B$, so they will tend to converge to their counterpart centroids in space $B$, rather than arbitrary cluster centroids. This way, the seeded clustering will only correct minor errors in the transformation, and move the centroids to their “right positions” in space $B$. Similar to the first strategy, we use exponential smoothing. [^2] Note that the above justification for Refine-2 is our a priori rationale for using this method, and we do not make any claims as to whether this is the true reason for the ability of this method to improve the alignment. Our only claim is that we observe this step improves the alignment consistently, and despite the fact that the Refine-1 has already converged and no longer shows improvement.
Crucially, we find that one iteration of Refine-2 improves the alignment obtained by Refine-1. Curiously, two or more lead to a slight deterioration of the alignment, although this might be partly explained by our use of a number of clusters smaller than the dimensionality.
5 Experiments
We replicate the in-distribution experiment from Jha et al. (2025). We run the algorithm for all pairs among four sets of text encoders used in Jha et al. (2025): stella, gtr, granite, and e5 (Zhang et al., 2025; Ni et al., 2021; Awasthy et al., 2025; Wang et al., 2024). Experiments use K-Means clustering from scikit-learn (Pedregosa et al., 2011), QAP is solved with the scipy (Virtanen et al., 2020) implementation of 2-OPT (Croes, 1958).
We take a random subset of the Natural Questions dataset (Kwiatkowski et al., 2019) of size $60,000$ and compute sentence embeddings under all possible encoders. We leave $8192$ sentences for evaluation. For each pair, we split the remaining embeddings equally between the source encoder (space $A$) and target encoder (space $B$), so there is no overlap between the source and target sentences. All experiments are run on a Colab notebook with a CPU runtime, and repeated three times for each pair to demonstrate consistency. We compare the results to the numbers reported in Jha et al. (2025).
Metrics.
We evaluate alignment quality using two primary metrics. The first is Top-1 Accuracy, the fraction of queries where the nearest neighbor in the target space is the correct match, defined by
$\text{Accuracy}=\frac{1}{N}\sum_{i=1}^{N}\delta(\text{NN}(\mathbf{W}\mathbf{a}_{i})=\mathbf{b}_{i}),$
where $\delta$ is the indicator function, $\text{NN}(\cdot)$ is the nearest neighbor of source embedding $\mathbf{a}_{i}$ in the target space, and $\mathbf{b}_{i}$ is its true corresponding target embedding. The second is the Average Rank, the mean position of the true match in the sorted list of distances from the query to all targets,
$\text{AvgRank}=\frac{1}{N}\sum_{i=1}^{N}\text{rank}(\mathbf{b}_{i}|\mathbf{a}_{i}).$
We also report the mean cosine similarity after each step of the algorithm to show its progression over the stages. [^3] Cosine similarity in itself is indicative, but it can also be misleading. Cosine similarity in mini-vec2vec underestimates the similarity, in the original space, due to the removal of the mean vector, which is a dominant part of all vectors, accounting for $\sim 30\%-70\%$ of the vector’s norm.
Hyperparameters.
We have found that the method is extremely robust to hyperparameter choices, as long as they are in a reasonable range. We use $s=30$ runs of relative representation alignment with $c=20$ clusters. We use $k=50$ neighbors in Mapping Estimation. We set the sampling rate in Refine-1 to $n_{s}=10,000$ and utilize $k^{\prime}=50$ neighbors; In Refine-2, we use $c^{\prime}=500$ clusters. For exponential smoothing, we use $\alpha=0.5$.
Note this is a rather economical choice of hyperparameters, which we would not have suggested to use a priori. It turned out to work well, regardless. The full hyperparameter configuration used in the experiments is also documented in the code linked.
Results.
In Table 1, we demonstrate that our approach matches or exceeds vec2vec across (almost) the entire table. Our advantage is more noticeable in the rank metric. Importantly, the method is extremely robust, as is evidenced by the small standard deviation, already at the end of the first refinement step. Unlike vec2vec, our method does not collapse on the $e5$ and gtrgtr pair.
We underline scores that are within a $0.01$ margin, as they might result from rounding noise and stochasticity, and a slightly different experimental setup. [^4] Our results are calculated over multiple runs and a single evaluation set, while vec2vec computes the results of a single run over 8 evaluation sets. Due to the small variance in most experiments, in both methods, we believe that a $0.01$ margin is a good compromise. Our results are not meant to show superiority, but rather matching performance for a fraction of the cost – except in the cases where vec2vec converges poorly, in which we observe mini-vec2vec’s superior behavior. Computationally, one run is completed on a CPU in less than ten minutes, while vec2vec requires 1–7 days on a GPU, depending on the hardware.
Analysis.
We observe that despite variation in the approximate matching phase (Initial), Refine-1 consistently improves cosine similarity and converges to stationary points with nearly identical cosine similarity values across different random initializations. This reproducible convergence behavior suggests the optimization landscape may contain well-connected regions or multiple basins with similar minima. The convergence of Refine-1 appears to reach a stationary point of the implicit nearest-neighbor matching objective. However, applying Refine-2 successfully improves upon this converged solution. The two algorithms appear to optimize different implicit proxy objectives, and both contribute to reaching a good solution.
6 Related Work
The study of universality has its roots in several works. Representation alignment is well-studied in the supervised setup. Lenc and Vedaldi (2015) and Bansal et al. (2021) demonstrated that linear transformations are often sufficient for aligning neural network representations through their stitching approach. Olah (2015) studied the equivalence of neural networks in relative terms. Moschella et al. (2023) have used these insights to encode a universal space where data points are represented relative to anchor points, utilizing it for zero-shot stitching. Huh et al. (2024) have aggregated and unified previous work, under the name Platonic Representation Hypothesis.
The task of usupervised alignment of embeddings has a long history in word embeddings (Conneau et al., 2018; Chen and Cardie, 2018; Grave et al., 2018) and machine translation (Lample et al., 2018; Artetxe et al., 2018). Most of the works rely on training a GAN. They often require a small seed of matched pairs in order to work. Hoshen and Wolf (2018) avoid the adversarial setup and use a three-step algorithm similar to ours. Their approach, specifically their approximate matching process, which relied on PCA, was insufficient for our purposes.
7 Conclusion
We have presented mini-vec2vec, a linear approach to unsupervised embedding alignment that achieves competitive performance while offering substantial advantages in computational efficiency, training stability, and interpretability. Our method demonstrates that adversarial training may be unnecessary for reliable embedding alignment, provided that structural correspondences can be estimated accurately.
The success of our approach provides strong empirical evidence for the universal geometry hypothesis and suggests that the geometric regularities in learned representations are more robust and accessible than previously assumed. By decomposing the alignment problem into classical optimization components, we achieve both theoretical clarity and practical efficiency.
The efficiency and accessibility of our method could democratize research on embedding alignment and cross-modal representation learning. By reducing computational barriers, we enable broader participation in this research area and facilitate deployment in resource-constrained environments. From a security perspective, efficient alignment methods raise important considerations about intellectual property protection in embedding spaces. If alignment can be performed quickly and reliably, it may become easier to leak data in the scenario described in Jha et al. (2025).