### v3

# mini-vec2vec: Scaling Universal Geometry Alignment with Linear Transformations

Abstract
We build upon vec2vec, a procedure designed to align embedding spaces without parallel data. While it usually finds a near-perfect alignment, it is expensive and unstable. We present mini-vec2vec, a simple and efficient alternative that is highly robust and incurs a much lower computational cost. Moreover, the transformation that mini-vec2vec learns is linear. The method consists of three main stages: a tentative matching of pseudo-parallel embedding vectors, transformation fitting, and iterative refinement. Our linear alternative exceeds the original instantiation of vec2vec by orders of magnitude in efficiency, while matching or exceeding its alignment performance. The method’s stability and interpretable steps facilitate scaling and unlock new opportunities for adoption in new domains and fields.
1 Introduction
Sentence embeddings capture rich semantic structure in high-dimensional vector spaces. They encode complex linguistic relationships and enable downstream tasks, such as semantic similarity and information retrieval. However, different architectures, training procedures, and even different models with the same architecture produce embedding spaces that cannot be compared directly, because they use a different coordinate system, a different “internal language”. The problem of aligning embedding spaces without parallel supervision – that is, without pairs of matching encodings of the same data – has emerged as a fundamental challenge with far-reaching implications. Successful unsupervised alignment has multiple implications across many domains: multilingual natural language processing (conneau2017word; zhang-etal-2017-adversarial), cross-modal alignment between text and image representations (maniparambil2024; schnaus2025), and privacy concerns (song2020informationleakage; text2vec).
The recent vec2vec method (jha2025vec2vec) has demonstrated that unsupervised alignment is indeed possible using CycleGAN (zhu2017cyclegan). Their approach is theoretically grounded in a principle called the Universal Geometry or the Platonic Representation Hypothesis (huh2024platonic), which posits that well-trained representations, regardless of their specific architectures or training procedures, tend to converge toward geometrically similar spaces, even if their coordinate systems differ. This means that there is a way to rotate the coordinate systems so they “understand” each other – and vec2vec does precisely that; it identifies a rotation that makes the coordinate systems aligned. While vec2vec has proven effective, adversarial training frameworks like CycleGAN bring inherent challenges that limit their practical applicability. Training them is computationally intensive, requires careful hyperparameter tuning, and is subject to training instabilities like mode collapse, oscillatory behavior, or failure to converge (saxena2023gan_survey; goodfellow2017nips2016). Moreover, they typically demand substantial computational resources, including GPU acceleration and large amounts of training data – factors that may prohibit their utilization by most researchers and users.
Our point of departure is still Universal Geometry, but we take a different approach. We are motivated by the remarkable success of supervised alignment approaches, particularly neural network stitching methods (lenc2014stitching; bansal2021revisiting), which have demonstrated that simple affine transformations often suffice to align representations. But unfortunately, stitching assumes access to pairs of matching representations. Motivated by the Platonic Representation Hypothesis, we hypothesize that we can nonetheless uncover this linear alignment without even needing adversarial training. Specifically, we hypothesize that the similarity in the geometric structure – the relative arrangement of semantic clusters, the preservation of neighborhood structures, and the overall geometric organization of concepts – should provide sufficient signal for alignment without parallel supervision and without need for adversarial methods. To this end, we utilize the Relative Representations framework of moschella2022relative, which allows us to operate in a universal space agnostic of arbitrary coordinate systems. We then use it as a stepping stone to produce a noisy alignment that we refine with further steps. [^1] In Section 6, we discuss a variant that sidesteps the use of relative representations, but we tested it far less.
mini-vec2vec is a simple and robust pipeline with competitive alignment performance offering several key advantages over adversarial approaches: computational efficiency (very fast and running on commodity CPU hardware), training stability (good performance and robustness to hyperparameters), interpretability (each step has a clear geometric interpretation), and sample efficiency (effective with substantially fewer training examples). Through comprehensive experiments, we demonstrate that this simple approach not only matches but often exceeds the performance of the more complex adversarial method while requiring orders of magnitude less computational resources.
2 Preliminaries
The hypothesis that neural networks converge to geometrically similar representations has been explored across multiple domains and scales. Early work by colah2015visualizing has revealed striking structural similarities in representations of different convolutional networks, [^2] Olah briefly refers to even earlier ideas (e.g., a concept called “meta-tSNE” by Hinton), but we failed to locate them. suggesting that the geometric organization of learned features might be more universal than previously assumed. The Platonic Representation Hypothesis, formally articulated by huh2024platonic, represents the most comprehensive theoretical framework for understanding Universal Geometry. The hypothesis posits that well-trained representations converge toward a shared geometric structure that reflects the underlying structure of the data domain. Having the same geometry implies that the distance between objects should be independent of the models themselves. For example, if one wants to compute the semantic similarity of “I love your dog” and “I love my cat” using cosine similarity between their embeddings, the answer should be roughly the same regardless of the underlying embedding model. Intuitively, this empirical observation stands to reason, as semantic similarity originates from the semantic meaning that comes from the data itself. Below, we discuss the framework of moschella2022relative that extends this observation into a powerful tool.
Relative Representations
moschella2022relative have operationalized the observations of colah2015visualizing and proposed to use relative representations as a mathematically principled framework for comparing latent spaces. While absolute coordinates vary arbitrarily across models, we know that relative information, such as similarity between two objects, is (approximately) agnostic to the choice of embedding model. They propose to use a shared set of anchor points that other data points will be compared to as a way to ground the position of data points in a model-agnostic way.
Denote the space of data points by $\mathcal{Z}$ (e.g., sentences, images, or any other modality) and the embedding function by $E:\mathcal{X}\rightarrow\mathbb{R}^{d}$. Given a set of anchor data points $\{\mathbf{a}_{i}\}_{i=1}^{k}\subseteq\mathcal{Z}$, the relative representation of $\mathbf{z}\in\mathcal{Z}$ under $E$ is:
$\mathbf{r}_{E}(\mathbf{z})=[f(E(\mathbf{z)},E(\mathbf{a}_{1}));f(E(\mathbf{z}),E(\mathbf{a}_{2}));\ldots;f(E(\mathbf{z}),E(\mathbf{a}_{k}))],$
where $f(\cdot,\cdot)$ is a similarity function (typically cosine similarity) and ; represents concatenation.
The power of relative representations is that, as shown by colah2015visualizing and moschella2022relative, for two encoders $E_{1},E_{2}$, the relative representations of the same data point is roughly the same, i.e., $\mathbf{r}_{E_{1}}(\mathbf{z})\approx\mathbf{r}_{E_{2}}(\mathbf{z})$. In other words, the function $\mathbf{r}_{E}:\mathcal{X}\rightarrow\mathbb{R}^{D}$ is itself an embedding that maps a data point into $k$-dimensional space in a way that is approximately independent of $E$ (for a sufficiently well-trained $E$) and only depends on the choice of anchor points.
3 Problem Formulation
We follow the setup laid out in vec2vec. We assume access to two pools of text embeddings, each from a different encoder, and our task is to align their embedding spaces without paired examples. Similar to vec2vec, we learn a parameterized mapping (neural network or linear transformation) rather than a one-to-one matching between data points. This approach offers several advantages:
  • Generalization: New embeddings can be translated without re-applying the algorithm
  • Robustness: Works when one-to-one point matching doesn’t exist, requiring only similar data distributions
  • Out-of-distribution capability: The parameterized nature can (potentially) enable generalization beyond the training distribution
The Universal Geometry hypothesis indicates that such a task can be possible: “good enough” models learn equivalent geometries – relative similarities between data points remain approximately the same regardless of the embedding model.
Let $\mathcal{X}_{A}=\{\mathbf{x}^{A}_{i}\}_{i=1}^{n_{A}}$ and $\mathcal{X}_{B}=\{\mathbf{x}^{B}_{j}\}_{j=1}^{n_{B}}$ denote two sets of $d$-dimensional embeddings, where $\mathbf{x}^{A}_{i},\mathbf{x}^{B}_{j}\in\mathbb{R}^{d}$. The sentences come from the same underlying distribution $\mathcal{D}$, but – unlike most other works – there is no overlap between the sentences themselves – only the distribution is shared. We organize these embeddings into matrices $\mathbf{X}_{A}\in\mathbb{R}^{n_{A}\times d}$ and $\mathbf{X}_{B}\in\mathbb{R}^{n_{B}\times d}$, where each row corresponds to a single embedding vector. The fundamental challenge is that while both spaces may contain similar geometric structures, they are related by an unknown transformation.
Our goal is to learn a transformation function $f:\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that $f(E_{A}(\mathbf{z}))\approx E_{B}(\mathbf{z})$ where $\mathbf{z}$ represents a datapoint $\mathbf{z}\in\mathcal{Z}$. More generally, the setup described in jha2025vec2vec allows for learning a pair of functions $f_{A},f_{B}$ sending embeddings to a third space, such that $f_{A}(E_{A}(\mathbf{z}))\approx f_{B}(E_{B}(\mathbf{z}))$. We will not take advantage of this flexibility in this paper. [^3] It also complicates the analysis of the results. For example, $f_{A}=f_{B}=\text{const}$ would be a trivial solution if we don’t impose additional constraints. Similarity then needs to be assessed more carefully.
4 Methodology
4.1 Motivation
In this work, we restrict our attention to linear transformations, i.e., $f(\mathbf{x})=\mathbf{W}\mathbf{x}$ where $\mathbf{W}\in\mathbb{R}^{d\times d}$. Moreover, we will constrain our search space to orthogonal transformations, i.e., matrices $\mathbf{W}$ that satisfy $\mathbf{W}\mathbf{W}^{T}=\mathbf{W}^{T}\mathbf{W}=\mathbf{I}$. This is motivated by the Universal Geometry hypothesis – we expect cosine similarities to be preserved across embedding spaces. Later on, we will relax this assumption and ultimately the matrix will be an exponentially-smoothed average of orthogonal matrices, enforcing orthogonality softly.
Our linear approach, mini-vec2vec, addresses key limitations of the current approach: GANs are notoriously unstable and hard to train, demand extensive computational resources and training time. They often require tedious hyperparameter tuning and sometimes multiple re-runs to converge to good solutions. On the other hand, we know that in supervised cases (model stitching), there exists an affine mapping between the latent spaces, suggesting the linear case is realizable – the question is how to find it efficiently. mini-vec2vec achieves that, and it has several appealing properties: it requires significantly less data than vec2vec (60k vs. 1 million samples), runs on a CPU rather than requiring extensive GPU resources, and provides more stable training dynamics.
Our method resembles the approach described in hoshen-wolf-2018-non for word embeddings, proceeding in three main stages: approximate matching, transformation fitting, and iterative refinement. This paradigm is very powerful, and we adopt a similar approach. Unfortunately, text embeddings are somewhat more complicated than word embeddings, partly because there is no small exhaustive set of objects (words) that need to be matched. Moreover, the very concept of non-overlapping datasets is somewhat contrived for word embeddings, precisely because the set of possible objects is closed and relatively small. Therefore, word embedding alignment is rarely considered in the non-overlapping datasets setup (even in multilingual matching, words mostly have a ground-truth translation). We find that applying the algorithm from hoshen-wolf-2018-non naively to our use case doesn’t work. The approximate matching stage, in particular, requires more delicate treatment in our case. The iterative refinement step is also improved by a slightly more subtle approach.
def mini_vec2vec(X_A, X_B):
X_A, X_B = normalize(X_A), normalize(X_B)
C = anchor_pairs(X_A, X_B)
W = procrustes(C)
W = refine_1(X_A, X_B, W)
W = refine_2(X_A, X_B, W)
return W
def anchor_pairs(X_A, X_B):
R_A_list, R_B_list = [], []
for i in range(num_runs):
C_A = k_means(X_A, num_clusters)
C_B = k_means(X_B, num_clusters)
S_A = similarity_function(C_A, C_A)
S_B = similarity_function(C_B, C_B)
pi = solve_QAP(S_A, S_B)
C_B_permuted = permute(C_B, pi)
R_A_list.append(similarity_function(X_A, C_A))
R_B_list.append(similarity_function(X_B, C_B_permuted))
R_A, R_B = concatenate(R_A_list), concatenate(R_B_list)
neighbors = knn(R_A, R_B, num_neighbors=k)
X_matched = average_nearest_neighbors(X_B, neighbors)
return X_A, X_matched
Listing 1: Pseudo-code for mini-vec2vec (refinement algorithms deferred to Listing LABEL:lst:refine)
We will now describe our algorithm in detail. Pseudo-code is provided in Listing LABEL:lst:full_algo.
4.2 Preprocessing
We first normalize the embeddings in both spaces. We center each set around its respective mean and normalize the embeddings to the unit hypersphere:
$\boldsymbol{\mu}_{A}$ $=\frac{1}{n_{A}}\mathbf{X}_{A}^{T}\mathbf{1}_{n_{A}},\quad\tilde{\mathbf{X}}_{A}=\mathbf{X}_{A}-\mathbf{1}_{n_{A}}\boldsymbol{\mu}_{A}^{T}$
$\hat{\mathbf{X}}_{A}[i,:]$ $=\frac{\tilde{\mathbf{X}}_{A}[i,:]}{\|\tilde{\mathbf{X}}_{A}[i,:]\|_{2}},\quad\text{similarly for space }B$
where $\boldsymbol{\mu}_{A}$ is the mean of space $A$, $\mathbf{1}_{n_{A}}$ is the vector of ones of length $n_{A}$, and $\|\cdot\|_{2}$ denotes the L2 norm.
4.3 Approximate Matching
To establish structural correspondences, we employ an anchor discovery procedure based on multiple runs of clustering and matching. We will first describe a single run, and then explain how we robustify it by ensembling multiple runs:
Centroid Matching
The reason we choose to align cluster centroids is that they roughly represent the central landmarks in the dataset. Since the data comes from a shared distribution $\mathcal{D}$, we expect clusters to be similar overall, and all that remains to do is to find the alignment between them. Of course, clusters are not very consistent even with exactly the same data, so the approach needs some tolerance to errors (hence the robustification step below).
We perform k-means clustering on both embedding spaces independently to obtain cluster centroids $\{\mathbf{c}^{A}_{j}\}$ and $\{\mathbf{c}^{B}_{j}\}$ where $j=1,\ldots,k$. We then compute similarity matrices between the centroids independently in each embedding space:
$S^{A}_{ij}$ $=\cos(\mathbf{c}^{A}_{i},\mathbf{c}^{A}_{j}),\quad S^{B}_{ij}=\cos(\mathbf{c}^{B}_{i},\mathbf{c}^{B}_{j})$
where $\mathbf{c}^{A}_{i}$ and $\mathbf{c}^{B}_{i}$ are the $i$-th cluster centroids in spaces $A$ and $B$ respectively.
We then find the optimal matching between cluster centroids from different embedding spaces by solving the Quadratic Assignment Problem (QAP):
$\pi^{*}=\arg\max_{\pi\in\Pi_{k}}\;\sum_{i=1}^{k}\sum_{j=1}^{k}S^{A}_{ij}\,S^{B}_{\pi(i)\,\pi(j)}$
where $\pi:\{1,2,...,k\}\rightarrow\{1,2,...,k\}$ is a permutation from the set $\Pi_{k}$ of all permutations on $k$ elements. Due to the small number of cluster centroids, this takes only a few seconds.
Relative Representations
Once we have identified an alignment between $A\text{-}$ and $B\text{-centroids}$, we use the aligned centroids as anchors. We represent each embedding through its relationships to anchor points. For embedding $\mathbf{x}^{A}_{i}$, we construct:
$\mathbf{r}^{A}_{i}=[\cos(\mathbf{x}^{A}_{i},\mathbf{c}^{A}_{1}),\cos(\mathbf{x}^{A}_{i},\mathbf{c}^{A}_{2}),\ldots,\cos(\mathbf{x}^{A}_{i},\mathbf{c}^{A}_{k})]$
Similarly, for $\mathbf{x}^{B}_{i}$, we compute (note that we need to re-arrange the indices with $\pi$):
[…]
Robustification via Ensembling
We improve robustness by repeating the process described above and concatenating relative representations from multiple runs:
$\mathbf{r}^{A}_{i}=[\mathbf{r}^{A}_{i,1};\mathbf{r}^{A}_{i,2};\ldots;\mathbf{r}^{A}_{i,s}]\in\mathbb{R}^{sk},\qquad\text{similarly for space }B$
This concatenation creates richer structural signatures that are more robust to noise that comes from bad alignments.
4.4 Mapping Estimation
We construct pseudo-parallel pairs by sending each element in space $A$ to the average of its $k$ nearest neighbors from space $B$ (based on similarity in relative space):
$\mathcal{C}=\left\{\left(\mathbf{x}^{A}_{i},\frac{1}{k}\hskip-3.00003pt\sum_{\tiny\hskip 2.94443pt{\mathbf{r}^{B}_{j}\in\mathcal{N}_{k}(\mathbf{r}^{A}_{i})}\hskip-13.25018pt}\hskip-3.00003pt\mathbf{x}^{B}_{j}\,\,\,\,\right):\,i=1,\ldots,n_{A}\right\}$
where $\mathcal{N}_{k}(\mathbf{r}^{A}_{i})$ denotes the $k$ nearest neighbors to $\mathbf{r}^{A}_{i}$ in the concatenated relative representation space. Note that the neighborhood information is obtained from the relative space (which is shared) and averaging takes place in the absolute space $B$. Optimal orthogonal transformation is obtained by Procrustes analysis: $\mathbf{W}^{*}=\mathbf{V}\mathbf{U}^{T}$ where $\mathbf{U}\boldsymbol{\Sigma}\mathbf{V}^{T}=\text{SVD}(\mathbf{A}^{T}\mathbf{B})$, and $\mathbf{A}$, $\mathbf{B}$ are matrices formed from the pseudo-parallel pairs in $\mathcal{C}$.
4.5 Iterative Refinement
The initial transformation provides a coarse alignment that we refine with two complementary strategies. Pseudo-code is provided in Listing LABEL:lst:refine.
Refine-1: Matching-based Refinement
This refinement strategy proceeds by transforming embeddings from space $A$ to $B$ using $\mathbf{W}$, averaging their nearest neighbors in space $B$ based on cosine similarity, and obtaining a new orthogonal transformation with Procrustes analysis – similar to the process described in the Mapping Estimation stage (Section 4.4), with the crucial difference that now the similarity is computed in the ambient space $B$ rather than relative space. The method is cheap and it runs for many iterations (50–100), using subsampling to boost computational efficiency further.
We use exponential smoothing for transformation updates:
$\mathbf{W}^{(t+1)}=(1-\alpha)\mathbf{W}^{(t)}+\alpha\mathbf{W}_{\text{new}}^{(t)}$
Refine-2: Clustering-based Refinement
This strategy aims to improve the large-scale matching of embeddings by matching cluster centroids. We begin by clustering the embeddings in space $A$. We then transform the $A$ cluster centroids, and cluster the $B$ embeddings with the transformed $A$ centroids as initialization. The use of cluster centroids creates more stable pseudo-parallel pairs based on cluster assignments rather than individual nearest neighbors. The use of the transformed centroids as initialization follows from the assumption that they cannot be far off from cluster centroids in space $B$, so they will tend to converge to their counterpart centroids in space $B$, rather than arbitrary cluster centroids. This way, the seeded clustering will only correct minor errors in the transformation, and move the centroids to their “right positions” in space $B$. Similar to the first strategy, we use exponential smoothing.
Crucially, we find that one iteration of Refine-2 improves the alignment obtained by Refine-1. Two or more lead to a slight deterioration of the alignment, although this might be partly explained by our use of a number of clusters smaller than the dimensionality.
def refine_1(X_A, X_B, W):
for t in range(num_iters):
X_sample = sample(X_A, subsampling_rate)
X_transformed = X_sample @ W
neighbors = knn(X_transformed, X_B, num_neighbors=k2)
X_sample_matched = average_nearest_neighbors(X_B, neighbors)
W_new = procrustes(X_sample, X_sample_matched)
W = (1 - alpha) * W + alpha * W_new
return W
def refine_2(X_A, X_B, W):
C_A = k_means(X_A, num_clusters2)
C_B = k_means(X_B, num_clusters2, init=C_A @ W)
W_new = procrustes(C_A, C_B)
return (1 - alpha) * W + alpha * W_new
Listing 2: Pseudo-code for the iterative refinement step
5 Experiments
We replicate the main experiment from jha2025vec2vec. We run our algorithm for all pairs of encoders used in jha2025vec2vec. Details about the encoders are provided in Table 2. Experiments use k-means clustering from scikit-learn (scikit-learn) and QAP is solved with the scipy (2020SciPy-NMeth) implementation of 2-OPT. We re-run the 2-OPT alignment algorithm 30 times and choose the permutation with the highest alignment score. This helps improve the robustness of the alignment, as a single 2-OPT is often not enough.
We take a random subset of the Natural Questions dataset (natural_questions) of size $60,000$ and compute sentence embeddings under all possible encoders. We leave $8192$ sentences for evaluation. For each pair, we split the remaining embeddings equally between the source encoder (space $A$) and target encoder (space $B$), so there is no overlap between the source and target sentences. All experiments are run on a Colab notebook with a CPU runtime and repeated three times for each pair to demonstrate consistency. We compare our results to the numbers reported in jha2025vec2vec.
Metrics
We evaluate alignment quality using two primary metrics. The first is Top-1 Accuracy, the fraction of queries where the nearest neighbor in the target space is the correct match, defined by:
$\text{Accuracy}=\frac{1}{N}\sum_{i=1}^{N}\delta(\text{NN}(\mathbf{W}\mathbf{a}_{i})=\mathbf{b}_{i}),$
where $\delta$ is the indicator function, $\text{NN}(\cdot)$ is the nearest neighbor of source embedding $\mathbf{a}_{i}$ in the target space, and $\mathbf{b}_{i}$ is its true corresponding target embedding. The second is the Average Rank, the mean position of the true match in the sorted list of distances from the query to all targets,
$\text{AvgRank}=\frac{1}{N}\sum_{i=1}^{N}\text{rank}(\mathbf{b}_{i}|\mathbf{a}_{i}).$
We also report the mean cosine similarity after each step of the algorithm to show its progression over the stages. Cosine similarity in itself can be used as an evaluation metric, but it can also be misleading. Cosine similarity in mini-vec2vec underestimates the similarity in the original space, due to the removal of the mean vector, which is a dominant part of all vectors, accounting for $\sim 30\%-70\%$ of the vector’s norm. Also, setups that move the vectors to a third space, like vec2vec, have an advantage as they can change the geometry. Our setup does not change the geometry (apart from aparametric translation and scaling in the preprocessing step).
Hyperparameters
We have found that the method is extremely robust to hyperparameter choices, as long as they are in a reasonable range. We make an economical choice of hyperparameter values due to the large number of runs, but we find the choice to be sufficiently robust still. We use $s=30$ runs of relative representation alignment with $c=20$ clusters. In Mapping Estimation, we use $k=50$ neighbors. We set the sampling rate in Refine-1 to $n_{s}=10,000$ and utilize $k^{\prime}=50$ neighbors; In Refine-2, we use $c^{\prime}=500$ clusters. For exponential smoothing, we use $\alpha=0.5$. The full hyperparameter configuration used in the experiments is also documented in the code linked. For the $\textit{gtr}\leftrightarrow\textit{gte}$ pair, we increase the number of clusters to $c=30$. We find that for this pair, the algorithm often converges to a sub-optimal solution in the $c=20$ regime.
vec2vec mini-vec2vec
$M_{1}$ $M_{2}$ Top-1 $\uparrow$ Rank $\downarrow$ Initial Refine-1 Refine-2 Top-1 $\uparrow$ Rank $\downarrow$
gtr 0.99 1.19 (0.1) 0.25 (0.02) 0.58 (0.00) 0.57 (0.00) 0.99 (0.00) 1.02 (0.00)
gran. stel. 0.98 1.05 (0.0) 0.38 (0.00) 0.59 (0.00) 0.60 (0.00) 0.99 (0.00) 1.03 (0.00)
e5 0.98 1.11 (0.0) 0.28 (0.00) 0.52 (0.00) 0.53 (0.00) 0.99 (0.00) 1.05 (0.00)
gte 0.95 1.18 (0.0) 0.35 (0.04) 0.54 (0.00) 0.56 (0.00) 0.98 (0.00) 1.05 (0.00)
gran. 0.99 1.02 (0.0) 0.31 (0.02) 0.58 (0.00) 0.57 (0.00) 1.00 (0.00) 1.01 (0.00)
gtr stel. 0.99 1.03 (0.0) 0.29 (0.00) 0.53 (0.00) 0.54 (0.00) 0.98 (0.00) 1.03 (0.00)
e5 0.84 2.88 (0.2) 0.22 (0.04) 0.48 (0.00) 0.49 (0.00) 0.98 (0.00) 1.06 (0.00)
$\text{gte}^{*}$ 0.93 2.31 (0.1) 0.29 (0.02) 0.50 (0.00) 0.51 (0.00) 0.98 (0.00) 1.05 (0.00)
gran. 0.95 1.22 (0.0) 0.31 (0.01) 0.57 (0.00) 0.58 (0.00) 0.99 (0.00) 1.03 (0.00)
gte $\text{gtr}^{*}$ 0.91 2.64 (0.1) 0.26 (0.02) 0.51 (0.00) 0.52 (0.00) 0.98 (0.00) 1.04 (0.00)
stella 1.00 1.00 (0.0) 0.42 (0.02) 0.70 (0.00) 0.71 (0.00) 1.00 (0.00) 1.00 (0.00)
e5 0.99 5.91 (0.5) 0.33 (0.05) 0.60 (0.00) 0.61 (0.00) 1.00 (0.00) 1.00 (0.00)
gran. 0.98 1.08 (0.0) 0.37 (0.02) 0.58 (0.00) 0.59 (0.00) 0.99 (0.00) 1.02 (0.00)
stel. gtr 1.00 1.10 (0.0) 0.28 (0.02) 0.54 (0.00) 0.54 (0.00) 0.96 (0.00) 1.10 (0.00)
e5 1.00 1.00 (0.0) 0.43 (0.00) 0.61 (0.00) 0.62 (0.00) 1.00 (0.00) 1.00 (0.00)
gte 1.00 1.00 (0.0) 0.42 (0.00) 0.69 (0.00) 0.69 (0.00) 1.00 (0.00) 1.00 (0.00)
gran. 0.99 2.20 (0.2) 0.30 (0.00) 0.54 (0.00) 0.56 (0.00) 0.99 (0.00) 1.04 (0.00)
e5 gtr 0.82 2.56 (0.0) 0.22 (0.03) 0.49 (0.00) 0.51 (0.00) 0.96 (0.00) 1.10 (0.00)
stel. 1.00 1.00 (0.0) 0.44 (0.01) 0.64 (0.00) 0.65 (0.00) 1.00 (0.00) 1.00 (0.00)
gte 1.00 1.00 (0.0) 0.33 (0.00) 0.61 (0.00) 0.62 (0.00) 1.00 (0.00) 1.00 (0.00)
Table 1: Results show mean $\pm$ standard deviation over 3 runs. Bold indicates best performance. Underline indicates nearly equally good performance ($\leq 0.01$ difference). Bold and underline combined indicate better performance, but only within a $0.01$ margin.
For the $gtr\leftrightarrow gte$ pair, we modify the number of clusters to $c=30$ in the approximate matching stage.
Results
In Table 1, we demonstrate that our approach matches or exceeds vec2vec across (almost) the entire table. The improvement is most noticeable in the rank metric. Importantly, the method is extremely robust as is evidenced by the small standard deviation, already by the end of the first refinement step. Unlike vec2vec, our method does not collapse on the $\textit{e5}\leftrightarrow\textit{gtr}$ pair or the $\textit{gtr}\leftrightarrow\textit{gte}$ pair. We underline scores that are within a $0.01$ margin, as they might result from rounding noise, stochasticity, and a slightly different experimental setup. [^4] Our results are calculated over multiple runs and a single evaluation set, while vec2vec computes the results of a single run over 8 evaluation sets. Due to the small variance in most experiments, in both methods, we believe that a $0.01$ margin is good enough.
Analysis
The slightly superior results of mini-vec2vec are very encouraging, but even in pairs where the results are slightly worse or match vec2vec, the main point is that it is achieved for a fraction of the cost. One run of mini-vec2vec is completed on a CPU in less than ten minutes, while vec2vec requires 1–7 days on a GPU, depending on the hardware. Another important point from the point of view of robustness is that even where the vec2vec results are relatively low, mini-vec2vec still performs well, close to perfect performance.
Analyzing the convergence of our algorithm, we observe that despite variation in the approximate matching step (Initial), Refine-1 consistently improves cosine similarity and converges to stationary points with nearly identical cosine similarity values across different random initializations. This reproducible convergence behavior suggests the optimization landscape is well-structured. The convergence of Refine-1 appears to reach a stationary point of the implicit nearest-neighbor matching objective. However, applying Refine-2 successfully improves upon this converged solution. The two methods appear to optimize different implicit proxy objectives, and both contribute to reaching a good solution.
Model Params Backbone Paper Year
gtr 110M T5 gtr 2021
e5 109M BERT e5 2022
gte 109M BERT gte 2023
stella 109M BERT stella 2023
granite 278M RoBERTa granite 2024
Table 2: Embedding models used in our experiments
6 Discussion
6.1 Choice of Hyperparameters
mini-vec2vec is stable under different choices of parameters, often converging to similar stationary points. Moreover, unlike many machine learning algorithms, most hyperparameters have a clear interpretation and often have a monotone behavior. Here we share trends that we have observed empirically in some setups, but a rigorous experimental analysis has not been made.
  • Number of concatenated representations: Increasing the number of runs in the anchor discovery step reduces the number of the failure modes. This is because noisy assignments are diluted, emphasizing the signal from true assignments. Note that even with very few runs, good alignments can be reached at random but with smaller probability.
  • Number of iterations: We find that the alignment is monotonically increasing as the number of iterations grows, unlike some algorithms where increasing information might cause degradation. This makes it a safe bet to increase the number of iterations. Across all cases examined, though, we found convergence to be quite fast. Almost without exception, 100 iterations were enough, by a large margin.
  • Subsampling rate: Subsampling allows us to speed up the algorithm quite substantially, but it can increase noise. The larger the sample size is, the closer this approximation is to the data. As a rule of thumb, we generally considered higher sampling rates to be better (at the cost of slower performance), but we haven’t tested this rigorously.
Other hyperparameters are less obvious. Number of clusters in the approximate matching step is likely good in moderation. While increasing it will give us more landmark points for alignment, it can theoretically lead to instabilities because the assignment of very granular centroids can be noisy. Increasing number of neighbors is often beneficial, but it eventually leads to diminishing returns, and might even deteriorate performance when neighborhoods become very large. At moderate scales, it helps to reduce noise, as averaging more neighbors leads to a smoother estimate of the true aligned position. Note that this is more relevant here than in word embeddings, where essentially every point has (more or less) one ground-truth translation, and it makes sense to rely on a single neighbor. Increasing the number of neighbors to very large numbers might lead to over-smoothing, which may harm performance. Finally, increasing $\alpha$ leads to higher cosine scores, but they don’t necessarily translate to better rank and top-1 scores. $\alpha$ is implicitly responsible for the distance of the matrix from being orthogonal. Additionally, increasing $\alpha$ makes convergence slightly slower.
In summary, unlike many deep learning setups, GANs in particular, the range of hyperparameters that lead to strong performance is generously wide. Each hyperparameter is interpretable, and one can plausibly guess in which direction to modify the parameter to improve effectiveness, if needed.
6.2 Failure Modes
We find that despite the many experiments we ran, ranging diverse hyperparameter configurations, failure modes were extremely rare. In fact, we didn’t experience any except on the $gte\leftrightarrow gtr$ pair, and modifying the number of clusters to $c=30$ solved this, though experimentation with more runs has shown that failure modes still exist in the gtr $\rightarrow$ gte direction, they are just rarer. Crucially, failure modes are not catastrophic. Failure modes are characterized by worse top-1 accuracy and average rank – but the decrease is not catastrophic, leading to $\sim 90\%$ top-1 accuracy and 5105-10 average rank, much better than any naive baseline.
6.3 Alternative Approaches
Through anecdotal experiments, we have found that it might be possible to sidestep the use of relative representations in the approximate matching step. Instead, we can perform Procrustes Analysis directly on the concatenated aligned centroids. It seems to do slightly worse in the approximate matching step, but it converges normally after. We emphasize, as noted above, that we haven’t tested this thoroughly to be certain it generalizes. For future reference, we call this variant mini-vec2vec without relative representations”, but catchier names can be used if meaning is clear.
7 Related Work
The study of universality has its roots in several works. Representation alignment is well-studied in the supervised setup. lenc2014stitching and bansal2021revisiting have demonstrated that linear transformations are often sufficient for aligning neural network representations through their stitching approach. colah2015visualizing studied the equivalence of neural networks in relative terms. moschella2022relative have used these insights to encode a universal space where data points are represented relative to anchor points, utilizing it for zero-shot stitching. Similarly, dar2023analyzing have demonstrated zero-shot stitching using a method equivalent to relative representations with token embeddings as anchors. huh2024platonic have aggregated and unified previous work, under the name Platonic Representation Hypothesis.
The task of unsupervised alignment of embeddings also has a long history in word embeddings (conneau2017word; chen-cardie-2018-unsupervised; grave2018unsupervisedalignmentembeddingswasserstein) and machine translation (lample2018unsupervisedmachinetranslationusing; artetxe2018unsupervisedneuralmachinetranslation). Most of the methods rely on training a GAN, and they often require a small seed of matched pairs in order to work. hoshen-wolf-2018-non avoid the adversarial setup and use a three-step algorithm similar to ours. Their approach, especially the approximate matching step, which relied on PCA, was insufficient for our purposes.
8 Conclusion
We have presented mini-vec2vec, a linear approach to unsupervised embedding alignment that achieves competitive performance while offering substantial advantages in computational efficiency, training stability, and interpretability. Our method demonstrates that adversarial training may be unnecessary for reliable embedding alignment, and identifying structural correspondences is possible in other ways. The success of our approach provides another strong empirical evidence for the Universal Geometry hypothesis and suggests that the geometric regularities in learned representations are more robust and accessible than previously assumed. By decomposing the alignment problem into classical optimization components, we achieve both theoretical clarity and practical efficiency.
The efficiency and accessibility of our method can help democratize research on embedding alignment and cross-modal representation learning. By reducing computational barriers, we enable broader participation in this research area and facilitate deployment in resource-constrained environments. From a security perspective, efficient alignment methods raise important considerations about data privacy and intellectual property protection in embedding spaces. If alignment can be performed quickly and reliably, it may become easier to leak data in the scenario described in jha2025vec2vec.
Acknowledgements
We thank John X. Morris for useful feedback. We thank Google for making CPU and GPU notebooks freely available.